Five data entry operators work at the data processing department of the Georgia Bank. Each day for 10 days, the number of defective records in a sample of 200 records typed by these operators has been noted, as follows:

Sample

1

2

3

4

5

6

7

8

9

10

No. of Defectives

19

22

28

12

14

8

25

10

16

6

  1. What SPC chart is needed here?
  1. Using z=3, establish upper and lower control limits for the control chart and judge if the process is in control.
Answer #1
Sample No. of defects p
1 19 0.095
2 22 0.11
3 28 0.14
4 12 0.06
5 14 0.07
6 8 0.04
7 25 0.125
8 10 0.05
9 16 0.08
10 6 0.03

pbar = sum of all p values / number of samples

=0.08

Here, p chart would be the most appropriate.

UCL = pbar + 3 [ pbar ( 1-pbar)/n]1/2 = 0.08 + 3 x [ 0.08x0.92/200]1/2 = 0.1375

LCL = pbar - 3 [ pbar ( 1-pbar)/n]1/2 = 0.08 - 3 x [ 0.08x0.92/200]1/2 = 0.0224

The process is out of control as one value is out of control limits.